joe is 1.59 m tall, wishes to find the height of a tree. He walks 23.56 m from the base of the tree along the shadow of the tree until his head is in a position where the tip of his shadow exactly overlaps the end of the tree top's shadow. He is now 8.42 m from the end of the shadows. How tall is the tree?
Answer by Theo(13342)

if I understand this correctly, the scenario is as follows:


           A   x  <--- top of tree
               x   x
               x      x
               x         x
               x            x
               x               x     D
               x                  x  
               x                     x   <--- top of joe's head
               x                     x   x
               x                     x      x
               x          1.59m -->  x         x
               x                     x            x
               x                     x               x
               x                     x                  x
           B   x   x   x   x   x   x   x   x   x   x   x   x   C
               |--------23.56m-------|--------8.42m--------|
                                     
                                     E



If this is the case, then we can first find the angle DCE and then find the height of the tree.

Tan (DCE) = opposite / adjacent = 1.59 / 8.42 = .188836105

Angle DCE = Arctan (.188836105) = 10.69359057 degrees.

Tan (ACB) = Tan (10.69359057) = .188836105

Tan (ACB) = opposite / adjacent = AB / BC = AB / (23.56 + 8.42) = AB / 31.98

Tan (ACB) = AB / 31.98

Since Tan (ACB) is the same as Tan (DCE), this formula becomes:

.188836105 = AB / 31.98

multiply both sides of this equation by 31.98 to get:

31.98 * .188836105 = AB

AB = 6.038978622 meters high.

Since triangle DCE is similar to triangle ACB, then the corresponding sides are proportional to each other so you could have used a ratio to find the same answer.

The ratio would have been:

DE / CE = AB / CB

This would have come out as:

1.59 / 8.42 = x / 31.98

Cross multiply to get 8.42*x = 1.59 * 31.98

divide both sides by 8.42 to get x = (1.59 * 31.98) / 8.42 = 6.038978622.